### v1

# Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime

Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime
Abdelmalek Boumali abdelmalek.boumali@univ-tebessa.dz or boumali.abdelmalek@gmail.com Laboratory of theoretical and applied Physics,
Echahid Cheikh Larbi Tebessi University, Algeria
(April 1, 2025)
This study explores the Feshbach-Villars (FV) formalism for spin-1/2 particles in curved spacetime. We derive the Hamiltonian form of the Dirac equation in this context and extend the FV transformation accordingly. The generalized Klein-Gordon equation, obtained by squaring the Dirac operator, is reformulated using the FV approach. We present the resulting Hamiltonian in both matrix and Pauli matrix forms, considering gravitational and electromagnetic interactions. The formalism is examined in both (1+2) and (1+3) dimensional spacetimes, with special attention given to the spin-field interaction term. This study provides a framework for studying relativistic quantum mechanics in curved spacetime, offering insights into the interplay between quantum effects, gravity, and electromagnetism.
I Introduction
Relativistic quantum mechanics represents a significant academic achievement that focuses on integrating the concepts of relativity and quantum mechanics. This integration expanded the framework of quantum theory and introduced the concept of negative energy levels, which emerged directly from relativistic formulations. Negative energy states are illustrated by the concept of the Dirac sea, which is an infinite expanse of particles with negative energy. In this model, the absence of a particle in the Dirac sea indicates the presence of a physical particle, specifically, a positron. This interpretation arises from the requirement that the energy of free particles must be positive, as dictated by the physical laws. By treating particles and antiparticles differently, this approach creates intrinsic asymmetry. While this asymmetry is essential for upholding the probabilistic principles of quantum physics, as outlined in the continuity equation, it also highlights the elegance of Dirac’s sea model. This concept is particularly applicable to fermions, which have a half-integer spin and obey the Pauli exclusion principle. However, this does not apply to bosons, which possess an integer spin. Quantum field theory (QFT) provides a more comprehensive framework by treating antiparticles as fundamental components of field interactions rather than relying on the Dirac sea [1, 2, 3].
In 1958, Feshbach and Villars [1] published their findings on linearizing the Klein-Gordon equation. Their work led to the development of the Feshbach-Villars (FV-0) equations, which allowed the Klein-Gordon equation to be expressed in the Schrödinger form, also known as the Hamiltonian form. This formalization addresses various contradictions in the Klein-Gordon equation, including the existence of negative energies. This facilitated a clearer interpretation of probability theory, particularly when considering particles with corresponding antiparticles.
Feshbach and Villars’ formalism resolved the key issues that had previously caused the abandonment of the Klein-Gordon theory, enabling a reevaluation of its principles. Significant research has been conducted using the Feshbach-Villars equations. Furthermore, efforts have been made to obtain numerical solutions and approximations for these equations, providing physicists with the means to tackle ongoing challenges and develop insights related to relativistic quantum mechanics (RQM) [4, 5, 6, 7, 8, 9, 10].
To expand the theory within the framework of general relativity, Silenko [11, 12] was the first to introduce a new form of conformal invariance for massless scalar particles in Riemannian spacetimes. He studied scalar particles in arbitrary spacetimes under general relativity (GR). To derive a Hamiltonian form of the covariant Klein-Gordon (KG) equation for both massive and massless particles, he utilized the generalized Feshbach-Villars transformation, as proposed in a previous study. Silenko then derived the general quantum-mechanical equations of motion and established their classical limits. Recently, this method has been applied in various contexts, and the results have been promising [13, 14, 15, 16].
The Feshbach-Villars equations have simplified the reevaluation of relativistic quantum mechanics for spin-0 particles and opened new avenues for investigating similar theories related to spin-1/2 particles. In addition, this formalism offers an alternative approach for spin-zero particles characterized by the Klein-Gordon equation (KG-0). This method incorporates antiparticles through charge density instead of probability density, thereby expanding the representation space to a pseudo-Hilbert space. Notably, this framework treats particles and antiparticles symmetrically, eliminating the need for a Dirac sea and enhancing the mathematical representation. This has significant implications for our understanding of the spin-zero particles and their interactions.
Building on this idea, FV-like equations have been proposed for spin-½ particles (KG–½). These equations have the potential to achieve complete particle-antiparticle symmetry and eliminate the need for a Dirac sea, representing a promising development. By altering the manner in which wave functions depict physical states, this extension alters certain scientific interpretations, particularly those involving transition probabilities. The differences become significant when examining phenomena where the influence of the Dirac sea is evident. The formalism presented by Feshbach and Villars paved the way for a new relativistic quantum mechanics equation for spin-1/2 particles. The worked developed by Robson and Staudte [2, 3] followed the idea of the Feshbach-Villars linearization procedure and applied it to the second-order Dirac equation. The result was an eight-component relativistic equation in Hamiltonian form, which preserves some of the inner product properties and the probabilistic interpretation obtained from the FV-0 equation. This eight-component equation or Feshbach-Villars equation for spin- 1 / 2 particles (FV-1/2) is relativistic and different from the Dirac equation.
This study aims to extend the Feshbach-Villars (FV) formalism for spin-1/2 particles to curved spacetime. Specifically, we analyze fermion particles in arbitrary spacetimes within the framework of general relativity (GR). To derive the Hamiltonian form of the covariant Dirac equation in curved spacetime, we employ the generalized Feshbach-Villars transformation introduced in Ref. [12]. It is worth noting that, to the best of our knowledge, there is a notable gap in the literature regarding this type of study.
II Hamiltonian form of Dirac equation in curved spacetime
The Hamiltonian form of the Dirac equation in curved spacetime involves recasting the covariant form of the equation into a time-evolution operator that acts on the wavefunction.
The Dirac equation in curved spacetime is:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\Psi=0,$ (1)
where:
  • $\gamma^{\mu}=e_{a}^{\mu}\gamma^{a}$are the curved spacetime gamma matrices, with $e_{a}^{\mu}$ being the tetrads and $\gamma^{a}$ the flat spacetime gamma matrices.
  • $\mathcal{D}_{\mu}=\partial_{\mu}+\frac{1}{4}\omega_{\mu}^{ab}\gamma_{ab}$ is the covariant derivative, including the spin connection $\omega_{\mu}^{ab}$.
  • $m$ is the mass of the spinor field.
Expanding, and rearranging to isolate the time derivative we have
$i\gamma^{0}\mathcal{D}_{0}\Psi=\left[-i\gamma^{i}\mathcal{D}_{i}+m\right]\Psi.$ (2)
wher
$\mathcal{D}_{0}=\partial_{0}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab},\,\mathcal{D}_{i}=\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}$ (3)
. (4)
Multiply (2) through by$(\gamma^{0})^{-1}$ (note: in most conventions, $(\gamma^{0})^{-1}=\gamma^{0}$), give
$i\partial_{0}\Psi=\gamma^{0}\left[-i\gamma^{i}\left(\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}\right)+m-\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}\right]\Psi.$ (5)
This gives the Hamiltonian form:
$i\partial_{0}\Psi=\mathcal{H}\Psi,$ (6)
where
$\mathcal{H}=\gamma^{0}\left[-i\gamma^{i}\mathcal{D}_{i}+m-\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}\right].$ (7)
Here
  • $-i\gamma^{0}\gamma^{i}\mathcal{D}_{i}$: Represents the kinetic energy and coupling of the spinor to the spatial curvature via the spin connection $\omega_{i}^{ab}$.
  • $m\gamma^{0}$: Represents the mass term in the Hamiltonian.
  • $-\frac{1}{4}\gamma^{0}\omega_{0}^{ab}\gamma_{ab}$: Represents the coupling to the spin connection in the time direction.
Final Form of the Hamiltonian
$\mathcal{H}=\gamma^{0}\left[-i\gamma^{i}\left(\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}\right)+m-\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}\right].$ (8)
This Hamiltonian governs the time evolution of the spinor field $\Psi$ in curved spacetime, accounting for both gravitational and gauge interactions.
The probability density for the Dirac equation in curved spacetime is a generalization of the flat spacetime case, reflecting the local geometry and the covariant nature of the wavefunction. In curved spacetime, the probability density is closely tied to the conserved current derived from the Dirac equation. These quantities is defined as the time component of the conserved current $J^{\mu}$, where:
$J^{\mu}=\bar{\psi}\gamma^{\mu}\psi,$ (9)
and $\bar{\psi}=\psi^{\dagger}\gamma^{0}$ is the Dirac adjoint. The conservation of the current $J^{\mu}$ is guaranteed by the Dirac equation:
$\nabla_{\mu}J^{\mu}=0,$ (10)
where $\nabla_{\mu}$ is the covariant derivative compatible with the curved spacetime metric. Thus, the probability density $\rho$ is the 00-component of the current:
$\rho=J^{0}=\bar{\psi}\gamma^{0}\psi.$ (11)
In curved spacetime, the interpretation of $\rho$ is influenced by the metric determinant $\sqrt{-g}$, which ensures proper normalization. To properly normalize the wavefunction in curved spacetime, we integrate $\rho$ over a spacelike hypersurface $\Sigma$:
$\int_{\Sigma}\rho\,\sqrt{-g}\,\text{d}^{3}x=1,$ (12)
where $\sqrt{-g}$ is the determinant of the spacetime metric $g_{\mu\nu}$, ensuring the proper volume element in curved spacetime. Despite the curvature, $J^{\mu}$ is conserved globally due to $\nabla_{\mu}J^{\mu}=0$, ensuring that the total probability remains constant.
III Feshbash-Villars of a particle of spin-1/2 in curved spacetime: The formalism
The modified Klein-Gordon equation obtained by squaring the operator in the Dirac equation, first found by Erwin Schrödinger is given by
$\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0$ (13)
where $R$ is the Ricci scalar, and $F_{\mu\nu}$ is the field strength of $A_{\mu}$ and
$\mathcal{D}_{\mu}=\partial_{\mu}+\frac{1}{4}\omega_{\mu}^{ab}\gamma_{ab}-iqA_{\mu}$ (14)
Following the procedure used by Silenko (12, 2025) the general form of the covariant KG equation reads(13) (see Annex A)
$\left(\partial_{0}^{2}+\frac{1}{g^{00}\sqrt{-g}}\left\{\partial_{i},\sqrt{-g}g^{0i}\right\}\partial_{0}+\frac{1}{g^{00}\sqrt{-g}}\partial_{i}\sqrt{-g}g^{ij}\partial_{j}+\frac{m^{2}-\lambda R}{g^{00}}\right)\psi=0.$ (15)
[…]
$\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)$ $=\mathcal{D}_{0}\left(\sqrt{-g}g^{00}\mathcal{D}_{0}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{01}\mathcal{D}_{1}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{02}\mathcal{D}_{2}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{03}\mathcal{D}_{3}\right)$
$+\mathcal{D}_{1}\left(\sqrt{-g}g^{10}\mathcal{D}_{0}\right)+\mathcal{D}_{1}\left(\sqrt{-g}g^{11}\mathcal{D}_{1}\right)+\mathcal{D}_{1}\left(\sqrt{-g}g^{12}\mathcal{D}_{2}\right)+\mathcal{D}_{1}\left(\sqrt{-g}g^{13}\mathcal{D}_{3}\right)$
$+\mathcal{D}_{2}\left(\sqrt{-g}g^{20}\mathcal{D}_{0}\right)+\mathcal{D}_{2}\left(\sqrt{-g}g^{21}\mathcal{D}_{1}\right)+\mathcal{D}_{2}\left(\sqrt{-g}g^{22}\mathcal{D}_{2}\right)+\mathcal{D}_{2}\left(\sqrt{-g}g^{23}\mathcal{D}_{3}\right)$
$+\mathcal{D}_{3}\left(\sqrt{-g}g^{30}\mathcal{D}_{0}\right)+\mathcal{D}_{3}\left(\sqrt{-g}g^{31}\mathcal{D}_{1}\right)+\mathcal{D}_{3}\left(\sqrt{-g}g^{32}\mathcal{D}_{2}\right)+\mathcal{D}_{3}\left(\sqrt{-g}g^{33}\mathcal{D}_{3}\right)$ (16)
So, divided by $g^{00}$ we obtain
$\left(D_{0}^{2}+\frac{1}{g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\}D_{0}+\frac{1}{g^{00}\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}+\frac{m^{2}-\frac{R}{4}}{g^{00}}\right)\psi=0.$ (17)
Which is a direct generalization of Eq. (15).
To obtain the Hamiltonian for the given generalized Klein-Gordon equation and its formulation, we proceed step by step using the Feshbach-Villars (FV) transformation and the framework provided in the question: The generalized Klein-Gordon equation in covariant form is given as:
$\left(D_{0}^{2}+\frac{1}{g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\}D_{0}+\frac{1}{g^{00}\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}+\frac{m^{2}-\frac{R}{4}}{g^{00}}\right)\psi=0,$ (18)
where $D_{\mu}=\partial_{\mu}+\Gamma_{\mu}$ represents the covariant derivative.
The FV transformation is applied to reformulate the wave function $\psi$ in terms of two components $\phi$ and $\chi$:
$\psi=\phi+\chi,\quad i\left(D_{0}+\mathrm{Y}\right)\psi=m(\phi-\chi),$ (19)
where:
$\mathrm{Y}=\frac{1}{2g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\},$ (20)
Using the FV transformation, the Klein-Gordon equation is rewritten as a first-order equation for $\psi$.
Substituting the definitions of $\phi$ and $\chi$, we derive a coupled system of equations:
$i\partial_{0}\psi=\mathcal{H}\psi,$ (21)
where the Hamiltonian operator $\mathcal{H}$ in terms of the components becomes:
$\mathcal{H}=\begin{pmatrix}\mathrm{Y}+\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}&-\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)\\ -\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)&-\mathrm{Y}-\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}\end{pmatrix}.$ (22)
For practical cases, simplifications often involve: (i) Assuming specific forms of the metric $g_{\mu\nu}$, (ii) Neglecting terms like $\frac{R}{4}$ or $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ in weak field approximations, and (iii) Approximating $\mathrm{Y}$ in terms of first-order derivatives of the metric and connections. The Hamiltonian is now expressed in a block-matrix form suitable for analysis of dynamics. The FV transformation splits $\psi$ into positive and negative energy states $(\phi,\chi)$.
To rewrite the Hamiltonian in terms of Pauli matrices, we use the Feshbach-Villars formalism and express the two-component wave function $\psi=(\phi,\chi)^{T}$. The generalized Klein-Gordon Hamiltonian takes the form:
$\mathcal{H}=\sigma_{z}\left(\mathrm{Y}+\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}\right)+\sigma_{x}\mathcal{O},$ (23)
where $\sigma_{x},\sigma_{y}$, and $\sigma_{z}$ are Pauli matrices, and the operators $\mathrm{Y}$ and $\mathcal{O}$ are defined as follows: Operator $\mathrm{Y}$:
$\mathrm{Y}=\frac{1}{2g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\},$ (24)
which includes contributions from the metric and the connections. Operator $\mathcal{O}$:
$\mathcal{O}=\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right),$ (25)
which accounts for the kinetic terms and electromagnetic interactions.
III.1 (1+2) dimension
The term $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ represents the interaction between the electromagnetic field tensor $F_{\mu\nu}$ and the spinor field via the spin matrix $\sigma^{\mu\nu}$. Here’s how to evaluate it: The field strength tensor is defined as:
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$ (26)
where $A_{\mu}$ is the electromagnetic four-potential. For a given electromagnetic configuration: $F_{0i}$: Electric field components $(E_{i})$. $F_{ij}$: Magnetic field components $(B_{k})$, with $\epsilon^{ijk}B_{k}=F_{ij}$.
In $(1+2)$-dimensional spacetime, $F_{\mu\nu}$ simplifies due to fewer spatial dimensions. The spin matrix for a spinor field is defined as:
$\sigma^{\mu\nu}=\frac{i}{2}\left[\gamma^{\mu},\gamma^{\nu}\right],$ (27)
where $\gamma^{\mu}$ are the gamma matrices satisfying the Clifford algebra:
$\{\gamma^{\mu},\gamma^{\nu}\}=2g^{\mu\nu}.$ (28)
In $(1+2)$-dimensional spacetime, the gamma matrices are typically:
$\gamma^{0}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\gamma^{1}=\begin{pmatrix}0&\sigma^{1}\\ \sigma^{1}&0\end{pmatrix},\quad\gamma^{2}=\begin{pmatrix}0&\sigma^{2}\\ \sigma^{2}&0\end{pmatrix},$ (29)
where $\sigma^{1}$ and $\sigma^{2}$ are the standard Pauli matrices. The commutator produces:
$\sigma^{\mu\nu}=\frac{i}{2}\begin{pmatrix}[\gamma^{\mu},\gamma^{\nu}]\end{pmatrix}.$ (30)
Using the decomposition of $F_{\mu\nu}$ into electric $(E_{i})$ and magnetic $(B_{i})$ fields:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}-e\vec{\Sigma}\cdot\vec{B},$ (31)
where: $\vec{E}$ is the electric field, $\vec{B}$ is the magnetic field and $\vec{\Sigma}$ is the spin operator.
For a purely magnetic field in $(1+2)$, with $F_{12}=B$, the spin interaction term reduces to:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=eB\sigma^{12}.$ (32)
If $\sigma^{12}$ corresponds to a specific representation of the spinor field in $(1+2)$-dimensional spacetime, this term captures the coupling of the spin to the magnetic field.
For a purely electric field in $(1+2)$, with $F_{01}=E_{x}andF_{02}=E_{y}$, the term becomes:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e(\gamma^{0}\gamma^{1}E_{x}+\gamma^{0}\gamma^{2}E_{y}).$ (33)
III.2 in (1+3)-dimensions
To evaluate $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ in $(1+3)$-dimensional spacetime, we need to compute the interaction between the electromagnetic field tensor $F_{\mu\nu}$ and the spin matrices $\sigma^{\mu\nu}.$
The electromagnetic field tensor in $(1+3)$-dimensions is:
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$ (34)
where $A_{\mu}=(A_{0},\vec{A})$ is the four-potential: Components of $F_{\mu\nu}$ are $F_{0i}=E_{i}$: Electric field components and $F_{ij}=\epsilon_{ijk}B_{k}$: Magnetic field components.
Explicitly:
$F_{\mu\nu}=\begin{pmatrix}0&-E_{x}&-E_{y}&-E_{z}\\ E_{x}&0&-B_{z}&B_{y}\\ E_{y}&B_{z}&0&-B_{x}\\ E_{z}&-B_{y}&B_{x}&0\end{pmatrix}.$ (35)
The spin matrix is:
$\sigma^{\mu\nu}=\frac{i}{2}\left[\gamma^{\mu},\gamma^{\nu}\right],$ (36)
where $\gamma^{\mu}$ are the gamma matrices satisfying the Clifford algebra:
$\{\gamma^{\mu},\gamma^{\nu}\}=2g^{\mu\nu}.$ (37)
In the Dirac representation:
$\gamma^{0}=\begin{pmatrix}I&0\\ 0&-I\end{pmatrix},\quad\gamma^{i}=\begin{pmatrix}0&\sigma^{i}\\ -\sigma^{i}&0\end{pmatrix},$ (38)
where $\sigma^{i}$ are the Pauli matrices. The commutators yield:
$\sigma^{\mu\nu}=\begin{cases}\sigma^{0i}=i\gamma^{0}\gamma^{i},&\text{time-space components,}\\ \sigma^{ij}=-\epsilon^{ijk}\begin{pmatrix}\sigma^{k}&0\\ 0&\sigma^{k}\end{pmatrix},&\text{space-space components.}\end{cases}$
Substituting $F_{\mu\nu}\text{and}\sigma^{\mu\nu}$:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}-e\vec{\Sigma}\cdot\vec{B}.$ (39)
With
$\vec{E}=(E_{x},E_{y},E_{z}),\quad\vec{\gamma}=(\gamma^{1},\gamma^{2},\gamma^{3}).$ (40)
The electric field term is:
$e\gamma^{0}\vec{\gamma}\cdot\vec{E}=e\gamma^{0}(\gamma^{1}E_{x}+\gamma^{2}E_{y}+\gamma^{3}E_{z}).$ (41)
Now, where
$\vec{B}=(B_{x},B_{y},B_{z}),\quad\vec{\Sigma}=\frac{i}{2}\epsilon_{ijk}\sigma^{ij}.$ (42)
The magnetic field term is:
$-e\vec{\Sigma}\cdot\vec{B}=-e\begin{pmatrix}\vec{\sigma}&0\\ 0&\vec{\sigma}\end{pmatrix}\cdot\vec{B}.$ (43)
Combining both terms:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}(\gamma^{1}E_{x}+\gamma^{2}E_{y}+\gamma^{3}E_{z})-e\begin{pmatrix}\vec{\sigma}\cdot\vec{B}&0\\ 0&\vec{\sigma}\cdot\vec{B}\end{pmatrix}.$ (44)
III.2.1 Special Case: Uniform Fields
  • Uniform Magnetic Field $(\vec{E}=0)$
    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=-e\vec{\Sigma}\cdot\vec{B}.$ (45)
  • Uniform Electric Field $(\vec{B}=0)$
    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}.$ (46)
IV Conclusion
The Feshbach-Villars formalism for a spin-1/2 particle in curved spacetime provides a robust method for exploring relativistic quantum mechanics in the presence of gravitational and electromagnetic fields. By deriving the Hamiltonian form of the Dirac equation and applying the FV transformation, we have developed a two-component formulation that clearly separates positive and negative energy states. This formalism presents several advantages:
  • It allows a more intuitive interpretation of the wavefunction components with respect to particle and antiparticle contributions.
  • The resulting Hamiltonian, expressed in terms of Pauli matrices, facilitates the analysis of spin-dependent effects in curved spacetime.
  • The approach naturally incorporates both gravitational and electromagnetic interactions, making it suitable for exploring various physical scenarios.
Our analysis in both (1+2) and (1+3) dimensions showcases the versatility of the formalism and highlights the specific forms of spin-field interactions across different spacetime configurations. The explicit treatment of the term involving the electromagnetic field tensor and spin matrices provides a clear understanding of how spin couples to external fields in curved spacetime. This work lays the groundwork for future studies on relativistic quantum effects in strong gravitational fields, the behavior of spin in curved spacetime, and potential applications in areas such as early universe cosmology and the physics of compact objects. Further research could investigate the implications of this formalism for quantum field theory in curved spacetime and explore possible experimental tests of the predicted effects.
?refname?
  • 1 (2025) H. Feshbach and F. M. H. Villars, Rev. Modern Phys. 30, 24 (1958).
  • 2 (2025) B. A. Robson and D. S. Staudte, J. Phys. A : Math. Gen 29, 157 (1996).
  • 3 (2025) D. S. Staudte, J. Phys. A 29, 169 (1996).
  • 4 (2025) M. Merad, L. Chetouani, and A. Bounames, Phys. Lett. A 267, 225 (2000).
  • 5 (2025) A. Bounames and L. Chetouani, Phys. Lett. A 279, 139 (2001).
  • 6 (2025) S. Haouat and L. Chetouani, Eur. Phys. J. C 41, 297 (2005).
  • 7 (2025) B Guettou1 and L Chetouani, Phys. Scr. 74, 12–18 (2006).
  • 8 (2025) N. Brown, Z. Papp, and R. M. Woodhouse, Few-Body Systems 57, 103 (2015).
  • 9 (2025) B. Motamedi, T. Shannon, and Z. Papp, Few-Body Systems 60, 65 (2019).
  • 10 (2025) D. Wingard, A. Garcia Vallejo and Z. Papp, Few-Body Syst 65, 30 (2024).
  • 11 (2025) A. J. Silenko, Phys. Rev. A 77, 012116 (2008).
  • 12 (2025) A. J. Silenko, Phys. Rev. D 88, 045004 (2013).
  • 13 (2025) A. Bouzenada, A. Boumali, E.O. Silva, Ann. Phys. 458, 169479 (2023)
  • 14 (2025) A. Bouzenada, A. Boumali, R. Vitória, F. Ahmed, M. Al-Raeei, Nucl. Phys. B 994, 116288 (2023)
  • 15 (2025) A. Bouzenada, A. Boumali, F. Ahmed, Nucl. Phys. B 1007, 116682 (2024)
  • 16 (2025) A.Bouzenada and A.Boumali. Ann. Physics 27, 169302 (2023).
  • 17 (2025) E. Schrodinger, Sitz. Preuss. Akad. Wiss. Berlin 105 (1932).
Appendix A Proof of Eq. (13)
The Dirac equation in curved spacetime is:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\Psi=0.$ (47)
Squaring the operator using:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\left[i\gamma^{\mu}\mathcal{D}_{\mu}+m\right],$ (48)
we get:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\left[i\gamma^{\mu}\mathcal{D}_{\mu}+m\right]=\left(i\gamma^{\mu}\mathcal{D}_{\mu}\right)^{2}-m^{2}.$ (49)
Expand $(i\gamma^{\mu}\mathcal{D}_{\mu})^{2}$:
$\left(i\gamma^{\mu}\mathcal{D}_{\mu}\right)^{2}=-\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (50)
The gamma matrices satisfy the Clifford algebra:
$\gamma^{\mu}\gamma^{\nu}=g^{\mu\nu}+\sigma^{\mu\nu},\quad\text{where }\sigma^{\mu\nu}=\frac{1}{2}\left[\gamma^{\mu},\gamma^{\nu}\right].$ (51)
Substitute:
$-\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=-g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}-\sigma^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (52)
The covariant derivative $\mathcal{D}_{\mu}=\partial_{\mu}+\Gamma_{\mu}+\text{spin connection}$ satisfies:
$[\mathcal{D}_{\mu},\mathcal{D}_{\nu}]=\frac{1}{4}R_{\mu\nu\lambda\sigma}\gamma^{\lambda\sigma}-iqF_{\mu\nu}.$ (53)
From this:
$g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R,$ (54)
where $R$ is the Ricci scalar.
For the spin terms:
$\sigma^{\mu\nu}F_{\mu\nu}=\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}.$ (55)
Substituting back:
$\left(i\gamma^{\mu}\mathcal{D}_{\mu}\right)^{2}=-\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}\right).$ (56)
Adding the mass term $-m^{2}$:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\left[i\gamma^{\mu}\mathcal{D}_{\mu}+m\right]\Psi=\left(-\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)+\frac{1}{4}R-\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0.$ (57)
The Klein-Gordon equation derived from the squared Dirac operator is:
$\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0.$ (58)
Here, the derivation is exact under the assumptions of: Minimal coupling to gravity and electromagnetism, Metric compatibility $(\nabla_{\mu}g_{\alpha\beta}=0)$ and Torsion-free spacetime. It is important to mention that (58) was originally derived by Schrodinger (17, 2025).

### v2

@@
 Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime
-Abdelmalek Boumali abdelmalek.boumali@univ-tebessa.dz or boumali.abdelmalek@gmail.com Laboratory of theoretical and applied Physics,
-Echahid Cheikh Larbi Tebessi University, Algeria
-(April 1, 2025)
-This study explores the Feshbach-Villars (FV) formalism for spin-1/2 particles in curved spacetime.
-We derive the Hamiltonian form of the Dirac equation in this context and extend the FV transformation accordingly.
-The generalized Klein-Gordon equation, obtained by squaring the Dirac operator, is reformulated using the FV approach.
-We present the resulting Hamiltonian in both matrix and Pauli matrix forms, considering gravitational and electromagnetic interactions.
-The formalism is examined in both (1+2) and (1+3) dimensional spacetimes, with special attention given to the spin-field interaction term.
-This study provides a framework for studying relativistic quantum mechanics in curved spacetime, offering insights into the interplay between quantum effects, gravity, and electromagnetism.
+Abdelmalek Boumali abdelmalek.boumali@univ-tebessa.dz or boumali.abdelmalek@gmail.com Echahid Cheikh Larbi Tebessi University, Tebessa, Algeria
+(February 24, 2026)
+This study extends the Feshbach–Villars (FV) formalism to spin–$1/2$ particles propagating in curved spacetime.
+Starting from the covariant Dirac equation, we derive its Hamiltonian form in a tetrad (vierbein) formulation and implement a diagonal FV transformation of the associated second-order (squared) equation.
+The resulting Schrödinger-type evolution equation is written in both matrix and Pauli-matrix representations and explicitly displays gravitational and electromagnetic couplings, including spin–field interaction terms.
+We analyze the formalism in $(1+2)$- and $(1+3)$-dimensional settings and apply it to cosmic-string backgrounds (static and spinning, with and without oscillator couplings), obtaining the corresponding spectra and wave functions.
+While the FV reformulation is algebraically equivalent to the covariant Dirac dynamics (up to the standard restriction that removes the spurious solutions introduced by squaring), it offers a complementary viewpoint that is useful in practice: (i) the particle/antiparticle sectors are represented by two coupled components with a simple (pseudo-)unitary inner product tied to the conserved charge; (ii) the effect of stationarity versus staticity is encoded directly in the FV operator $\mathrm{Y}$, which is controlled by the ADM shift vector and therefore diagnoses how frame–dragging backgrounds modify particle–antiparticle mixing; and (iii) the FV block structure provides a compact route to spectral problems in topologically nontrivial geometries.
 I Introduction
@@
 By treating particles and antiparticles differently, this approach creates intrinsic asymmetry.
-While this asymmetry is essential for upholding the probabilistic principles of quantum physics, as outlined in the continuity equation, it also highlights the elegance of Dirac’s sea model.
-This concept is particularly applicable to fermions, which have a half-integer spin and obey the Pauli exclusion principle.
-However, this does not apply to bosons, which possess an integer spin.
-Quantum field theory (QFT) provides a more comprehensive framework by treating antiparticles as fundamental components of field interactions rather than relying on the Dirac sea [1, 2, 3].
-In 1958, Feshbach and Villars [1] published their findings on linearizing the Klein-Gordon equation.
+Although this asymmetry is essential to maintain the probabilistic principles of quantum physics, as described in the continuity equation, it also highlights the elegance of Dirac’s sea model.
+This concept is particularly applicable to fermions that have a half-integer spin and obey the Pauli exclusion principle.
+However, this does not apply to bosons that possess an integer spin.
+Quantum field theory (QFT) provides a more comprehensive framework by treating antiparticles as fundamental components of field interactions rather than relying on the Dirac sea Feshbach and Villars (1958); Robson and Staudte (1996); Staudte (1996) .
+In 1958, Feshbach and Villars Feshbach and Villars (1958) published their findings on linearizing the Klein-Gordon equation.
 Their work led to the development of the Feshbach-Villars (FV-0) equations, which allowed the Klein-Gordon equation to be expressed in the Schrödinger form, also known as the Hamiltonian form.
@@
 Significant research has been conducted using the Feshbach-Villars equations.
-Furthermore, efforts have been made to obtain numerical solutions and approximations for these equations, providing physicists with the means to tackle ongoing challenges and develop insights related to relativistic quantum mechanics (RQM) [4, 5, 6, 7, 8, 9, 10].
-To expand the theory within the framework of general relativity, Silenko [11, 12] was the first to introduce a new form of conformal invariance for massless scalar particles in Riemannian spacetimes.
+Furthermore, efforts have been made to obtain numerical solutions and approximations for these equations, providing physicists with the means to tackle ongoing challenges and develop insights related to relativistic quantum mechanics (RQM) Merad et al.
+(2000); Bounames and Chetouani (2001); Haouat and Chetouani (2005); Guettoul and Chetouani (2006); Brown et al.
+(2015); Motamedi et al.
+(2019); Wingard et al.
+(2024).
+To extend the FV programme to general relativistic settings, Silenko Silenko (2008, 2013) identified a conformal invariance of the massless scalar theory in Riemannian spacetimes, i.e.
+an invariance of the properly defined wave equation under local Weyl rescalings $g_{\mu\nu}\to\Omega^{2}(x)g_{\mu\nu}$ (with a corresponding field rescaling).
+This property is useful because it constrains the curvature-coupling term and allows one to construct a Hamiltonian form that behaves consistently under conformal transformations.
 He studied scalar particles in arbitrary spacetimes under general relativity (GR).
@@
 Silenko then derived the general quantum-mechanical equations of motion and established their classical limits.
-Recently, this method has been applied in various contexts, and the results have been promising [13, 14, 15, 16].
+Recently, this method has been applied in various contexts, and the results have been promising Bouzenada et al.
+(2023a); Garah and Boumali (2025b, a); Bouzenada et al.
+(2023b); Bouzenada and Boumali (2023); Bouzenada et al.
+(2024).
 The Feshbach-Villars equations have simplified the reevaluation of relativistic quantum mechanics for spin-0 particles and opened new avenues for investigating similar theories related to spin-1/2 particles.
@@
 These equations have the potential to achieve complete particle-antiparticle symmetry and eliminate the need for a Dirac sea, representing a promising development.
-By altering the manner in which wave functions depict physical states, this extension alters certain scientific interpretations, particularly those involving transition probabilities.
+By recasting the second-order Dirac dynamics into an enlarged FV state space, the wave function is represented by coupled components associated with particle and antiparticle sectors.
+This change modifies the underlying inner product (from a probability-density norm to a charge-density norm), and therefore affects how one interprets amplitudes and transition probabilities in the single-particle theory.
 The differences become significant when examining phenomena where the influence of the Dirac sea is evident.
 The formalism presented by Feshbach and Villars paved the way for a new relativistic quantum mechanics equation for spin-1/2 particles.
-The worked developed by Robson and Staudte [2, 3] followed the idea of the Feshbach-Villars linearization procedure and applied it to the second-order Dirac equation.
+The worked developed by Robson and Staudte Robson and Staudte (1996); Staudte (1996) followed the idea of the Feshbach-Villars linearization procedure and applied it to the second-order Dirac equation.
 The result was an eight-component relativistic equation in Hamiltonian form, which preserves some of the inner product properties and the probabilistic interpretation obtained from the FV-0 equation.
@@
 To derive the Hamiltonian form of the covariant Dirac equation in curved spacetime, we employ the generalized Feshbach-Villars transformation introduced in Ref.
-[12].
+Silenko (2013).
 It is worth noting that, to the best of our knowledge, there is a notable gap in the literature regarding this type of study.
-II Hamiltonian form of Dirac equation in curved spacetime
+It is important to stress the scope of the present work in light of the well-known fact that the FV construction starts from (and therefore remains anchored to) the covariant Dirac theory.
+The motivation is not to propose a new relativistic wave equation, but to provide a Hamiltonian formulation in an enlarged FV state space in which (a) the separation into positive-/negative-frequency sectors becomes explicit, (b) the conserved inner product is naturally expressed in terms of a simple metric operator, and (c) the role of geometric data (staticity, shift vector, and topological identifications) in the mixing between sectors can be tracked directly at the operator level.
+These points are particularly transparent in stationary defect geometries with nontrivial $g_{0\varphi}$ or $g_{z\varphi}$ components, where the FV operator $\mathrm{Y}$ does not vanish and the azimuthal quantum number is shifted by energy- and momentum-dependent terms (see Sec.
+V).
+The paper is organized as follows.
+Section II reviews the FV linearization for the second-order Dirac equation in flat spacetime and establishes the notation used later.
+Section III derives the Hamiltonian form of the covariant Dirac equation in curved spacetime, emphasizing the tetrad fields and the spin connection.
+Section IV develops the FV formulation for spin–$1/2$ particles in general curved backgrounds and discusses the transformation properties of the resulting Hamiltonian.
+Section V applies the formalism to cosmic-string geometries (including oscillator couplings) and extracts the corresponding spectra.
+We conclude with a discussion of the scope and limitations of the approach and possible extensions.
+II Solution of Relativistic oscillators of Feshbach-Villars of Spin-1/2
+II.1 Theoretical framework
+The Dirac equation for a charged particle in an electromagnetic field is Greiner (2001):
+$\left(i\gamma^{\mu}D_{\mu}-m\right)\psi=0$ (1)
+where $D_{\mu}=\partial_{\mu}+\frac{iq}{\hbar c}A_{\mu}$, and $A_{\mu}$ is the electromagnetic four-potential. Multiply both sides by $\left(i\gamma^{\nu}D_{\nu}+mc\right)$:
+$\left(i\gamma^{\nu}D_{\nu}+mc\right)\left(i\gamma^{\mu}D_{\mu}-m\right)\psi=0$ (2)
+Expanding this gives:
+$\left[(i\gamma^{\nu}D_{\nu})(i\gamma^{\mu}D_{\mu})-m^{2}\right]\psi=0$ (3)
+or
+$\gamma^{\nu}\gamma^{\mu}D_{\nu}D_{\mu}\psi-m^{2}\psi=0$ (4)
+The gamma matrices satisfy the anticommutation relation:
+$\{\gamma^{\mu},\gamma^{\nu}\}=2g^{\mu\nu}$ (5)
+So,
+$\gamma^{\nu}\gamma^{\mu}=g^{\nu\mu}+\sigma^{\nu\mu}$ (6)
+where
+$\sigma^{\nu\mu}=\frac{1}{2}[\gamma^{\nu},\gamma^{\mu}]$ (7)
+Therefore,
+$\gamma^{\nu}\gamma^{\mu}D_{\nu}D_{\mu}=-g^{\nu\mu}D_{\nu}D_{\mu}+-\sigma^{\nu\mu}D_{\nu}D_{\mu}$ (8)
+The first term gives the d’Alembertian (wave operator), the second term gives the coupling to the electromagnetic field tensor $F_{\mu\nu}$ via the commutator:
+$D_{\nu},D_{\mu}]=iqF_{\nu\mu},\,F_{\nu\mu}=\partial_{\nu}A_{\mu}-\partial_{\mu}A_{\nu}$ (9)
+So,
+$\sigma^{\nu\mu}D_{\nu}D_{\mu}=q\sigma^{\nu\mu}F_{\nu\mu}$ (10)
+Putting it all together:
+$\left[D^{\mu}D_{\mu}-m^{2}+q\sigma^{\nu\mu}F_{\nu\mu}\right]\psi=0$ (11)
+The operator $D^{\mu}D_{\mu}$ is the covariant d’Alembertian, and $\sigma^{\nu\mu}F_{\nu\mu}$ gives the coupling to the electromagnetic field, which can be written in terms of electric and magnetic fields using the Pauli matrices and their combinations.
+The Dirac equation for a charged spin-1/2 particle in an electromagnetic field is "squared" to obtain a second-order equation (Eq. (5) in the reference Brown et al. (2015); Wingard et al. (2024)):
+$\left[D_{1}^{2}+D_{2}^{2}+D_{3}^{2}+D_{4}^{2}-\frac{m^{2}c^{2}}{\hbar^{2}}\right]\psi+\frac{q}{\hbar c}(\mathbf{B}\cdot\boldsymbol{\sigma}-i\mathbf{E}\cdot\boldsymbol{\sigma})\psi=0$ (12)
+where $D_{\mu}=\partial_{\mu}-\frac{iq}{\hbar c}A_{\mu}$, $\mathbf{B}=\nabla\times\mathbf{A}$, $\mathbf{E}=-\frac{1}{c}\frac{\partial\mathbf{A}}{\partial t}-\nabla\Phi$, and $\boldsymbol{\sigma}$ are the Pauli matrices.
+II.2 Feshbach-Villars Linearization
+The FV formalism rewrites the second-order equation as a first-order system in time by splitting the wave function :
+$\psi=\phi+\chi$ (13)
+$(i\frac{\partial}{\partial t}-V)\psi=m(\phi-\chi)$ (14)
+This leads to two coupled equations:
+$(i\frac{\partial}{\partial t}-V)(\phi+\chi)=m(\phi-\chi)$ (15)
+$(i\frac{\partial}{\partial t}-V)(\phi-\chi)=\frac{p^{2}}{m}(\phi+\chi)+m(\phi+\chi)-\frac{i}{m}(\nabla V\cdot\boldsymbol{\sigma})(\phi+\chi)$ (16)
+where $V=q\Phi$ is the scalar potential. These can be combined into a two-component Schrödinger-like equation:
+$i\frac{\partial}{\partial t}\begin{pmatrix}\phi\\ \chi\end{pmatrix}=H\begin{pmatrix}\phi\\ \chi\end{pmatrix}$ (17)
+with the FV Hamiltonian (Eq. (12) in the reference Wingard et al. (2024); Brown et al. (2015) ):
+$H=\begin{pmatrix}1&1\\ -1&-1\end{pmatrix}\frac{p^{2}}{2m}+\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}m+\begin{pmatrix}1&0\\ 0&1\end{pmatrix}V-\begin{pmatrix}1&1\\ -1&-1\end{pmatrix}\frac{i}{2m}(\nabla V\cdot\boldsymbol{\sigma})$ (18)
+Using Pauli matrices $\tau_{1},\tau_{2},\tau_{3}$ in the FV (component) space, the Hamiltonian is written as :
+$H=(\tau_{3}+i\tau_{2})\frac{p^{2}}{2m}+\tau_{3}m+I_{2}V-(\tau_{3}+i\tau_{2})\frac{i}{2m}(\nabla V\cdot\boldsymbol{\sigma})$ (19)
+[…]
+II.3 FV oscillator for particles of spin-1/2
+For the next discussion, to look at the DO in the FV representation, we begin with the minimal substitution $p\longrightarrow p-im\omega x$ of the momentum operator.
+Here $\omega$ is the oscillation frequency and $x_{\mu}=(0,x,0,0)$.
+Thus, by generalizing the momentum operator, the one dimensional FV Hamiltonian (19) becomes (Here $V=0$) Bouzenada et al.
+(2023a)
+$H=(\tau_{3}+i\tau_{2})\frac{\left(p+im\omega x\right)\left(p-im\omega x\right)}{2m}+\tau_{3}m$ (20)
+In this way, the Schrödinger formulation of the DO in one dimension is obtained by replacing the ansatz (13) into (15) with the following :
+$\frac{1}{2m}\left[\frac{d^{2}}{dx^{2}}-m^{2}\omega^{2}x^{2}+m\omega\right]\left(\phi+\chi\right)+m\phi$ $=E\phi$ (21)
+$-\frac{1}{2m}\left[\frac{d^{2}}{dx^{2}}-m^{2}\omega^{2}x^{2}+m\omega\right]\left(\phi+\chi\right)-m\chi$ $=E\chi$
+After making a calculation similar to that which was made in the previous equation of the FVO, it follows that
+$\left[\frac{d^{2}}{dx^{2}}-m^{2}\omega^{2}x^{2}+\left(E^{2}-m^{2}+m\omega\right)\right]\psi(x)=0$ (22)
+or,
+$\left[\frac{d^{2}}{dx^{2}}-\lambda_{1}x^{2}+\lambda_{2}\right]\psi(x)=0$ (23)
+where we have set,
+$\lambda_{1}=m^{2}\omega^{2},\quad\lambda_{2}=E^{2}-m^{2}+m\omega.$ (24)
+Eq.
+(23) is a second order differential equation for the field $\psi$ describing the FVO dynamics in one dimensional Minkowski space-time.
+The solution of the above equation can be found in the literature, and a quantification condition of energy results from this solution, providing.
+$\frac{E^{2}-m^{2}+m\omega}{2m}=\left(n+\frac{1}{2}\right)\omega,\quad n=0,1,2,\ldots$ (25)
+Having arranged and simplified the condition (24), we find the following expression of the energy spectrum
+$E=\pm\sqrt{2m\omega n+m^{2}}$ (26)
+Eq.
+(26) shows the relativistic energy spectrum of the FVO in space-time at Minkowski, which is the same as the one obtained for the KGO in flat space-time (see Ref.
+Boumali (2015)).
+We see that this result is expected because the energy spectrum is observable and is not affected by the representation we used.
+III Hamiltonian form of Dirac equation in curved spacetime
 The Hamiltonian form of the Dirac equation in curved spacetime involves recasting the covariant form of the equation into a time-evolution operator that acts on the wavefunction.
@@
   • $m$ is the mass of the spinor field.
+The tetrad (vierbein) fields $e^{a}{}_{\mu}$ and their inverse $e_{a}{}^{\mu}$ relate the curved metric to the local Minkowski metric via $g_{\mu\nu}=e^{a}{}_{\mu}e^{b}{}_{\nu}\,\eta_{ab}$ and define the curved gamma matrices $\gamma^{\mu}=e_{a}{}^{\mu}\gamma^{a}$.
+We distinguish: (i) the spacetime covariant derivative $\nabla_{\mu}$ acting on tensor indices, (ii) the gauge-covariant derivative $D_{\mu}=\partial_{\mu}-iqA_{\mu}$ used for minimal electromagnetic coupling, and (iii) the spinor covariant derivative $\mathcal{D}_{\mu}=\partial_{\mu}+\frac{1}{4}\omega_{\mu}{}^{ab}\gamma_{ab}-iqA_{\mu}$, which adds the spin connection to ensure local Lorentz covariance.
+These derivatives coincide on scalars but differ on tensors/spinors through their respective connection terms Nakahara (2003).
 Expanding, and rearranging to isolate the time derivative we have
 $i\gamma^{0}\mathcal{D}_{0}\Psi=\left[-i\gamma^{i}\mathcal{D}_{i}+m\right]\Psi.$ (28)
-wher
+where
 $\mathcal{D}_{0}=\partial_{0}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab},\,\mathcal{D}_{i}=\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}$ (29)
-. (4)
-Multiply (2) through by$(\gamma^{0})^{-1}$ (note: in most conventions, $(\gamma^{0})^{-1}=\gamma^{0}$), give
+$\gamma_{ab}=\frac{1}{2}[\gamma_{a},\gamma_{b}],$ (30)
+[…]
 $i\partial_{0}\Psi=\gamma^{0}\left[-i\gamma^{i}\left(\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}\right)+m-\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}\right]\Psi.$ (31)
@@
 Here
-  • $-i\gamma^{0}\gamma^{i}\mathcal{D}_{i}$: Represents the kinetic energy and coupling of the spinor to the spatial curvature via the spin connection $\omega_{i}^{ab}$.
+  • $i\gamma^{0}\gamma^{i}\mathcal{D}_{i}$: Represents the kinetic energy and coupling of the spinor to the spatial curvature via the spin connection $\omega_{i}^{ab}$.
   • $m\gamma^{0}$: Represents the mass term in the Hamiltonian.
-  • $-\frac{1}{4}\gamma^{0}\omega_{0}^{ab}\gamma_{ab}$: Represents the coupling to the spin connection in the time direction.
+  • $\frac{1}{4}\gamma^{0}\omega_{0}^{ab}\gamma_{ab}$: Represents the coupling to the spin connection in the time direction.
 Final Form of the Hamiltonian
@@
 $\nabla_{\mu}J^{\mu}=0,$ (36)
-where $\nabla_{\mu}$ is the covariant derivative compatible with the curved spacetime metric. Thus, the probability density $\rho$ is the 00-component of the current:
+where $\nabla_{\mu}$ is the covariant derivative compatible with the curved spacetime metric. Thus, the probability density $\rho$ is the 0-component of the current:
 $\rho=J^{0}=\bar{\psi}\gamma^{0}\psi.$ (37)
@@
 Despite the curvature, $J^{\mu}$ is conserved globally due to $\nabla_{\mu}J^{\mu}=0$, ensuring that the total probability remains constant.
-III Feshbash-Villars of a particle of spin-1/2 in curved spacetime: The formalism
-The modified Klein-Gordon equation obtained by squaring the operator in the Dirac equation, first found by Erwin Schrödinger is given by
+IV Feshbach-Villars of a particle of spin-1/2 in curved spacetime: The formalism
+The modified Klein-Gordon equation obtained by squaring the operator in the Dirac equation, first found by Erwin Schr"odinger Schrodinger (1932) is given by
 $\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0$ (39)
@@
 $\mathcal{D}_{\mu}=\partial_{\mu}+\frac{1}{4}\omega_{\mu}^{ab}\gamma_{ab}-iqA_{\mu}$ (40)
-Following the procedure used by Silenko (12, 2025) the general form of the covariant KG equation reads(13) (see Annex A)
+Following the procedure used by Silenko Silenko (2013) the general form of the covariant KG equation reads(39) (see Annex A)
 $\left(\partial_{0}^{2}+\frac{1}{g^{00}\sqrt{-g}}\left\{\partial_{i},\sqrt{-g}g^{0i}\right\}\partial_{0}+\frac{1}{g^{00}\sqrt{-g}}\partial_{i}\sqrt{-g}g^{ij}\partial_{j}+\frac{m^{2}-\lambda R}{g^{00}}\right)\psi=0.$ (41)
-[…]
+The curly bracket $\{\ldots,\ldots\}$ denotes the anticommutator. In our case, the first term is
 $\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)$ $=\mathcal{D}_{0}\left(\sqrt{-g}g^{00}\mathcal{D}_{0}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{01}\mathcal{D}_{1}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{02}\mathcal{D}_{2}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{03}\mathcal{D}_{3}\right)$
@@
 $\mathrm{Y}=\frac{1}{2g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\},$ (46)
-Using the FV transformation, the Klein-Gordon equation is rewritten as a first-order equation for $\psi$.
+The Feshbach-Villars Metric and Inner Product
+The FV transformation defined in Eq.
+(45) recasts the squared Dirac equation (Schr"odinger form) in Eq.
+(39) into a first-order Schr"odinger-type evolution equation in an enlarged state space.
+In particular, one introduces the two-component FV field
+$\Psi_{FV}=(\phi,\chi)^{T}$
+For spin–$1/2$, each component $\phi$ and $\chi$ is itself a Dirac spinor (four components in $3+1$ dimensions), so that $\Psi_{FV}$ carries eight components in total.
+Unlike the standard Dirac theory, the Feshbach-Villars transformation is generally not unitary with respect to the standard positive-definite $L^{2}$ inner product.
+Instead, it is unitary with respect to an indefinite metric operator associated with the conserved charge of the second-order theory.
+In the canonical FV basis, this metric operator is defined as $\eta_{FV}=\sigma_{3}\otimes\mathbb{I}_{4}$ (and reduces to $\eta_{FV}=\sigma_{3}$ in the scalar FV-0 case):
+$\eta_{FV}=\sigma_{3}\otimes\mathbb{I}_{4}=\begin{pmatrix}\mathbb{I}_{4}&0\\ 0&-\mathbb{I}_{4}\end{pmatrix}$.
+The corresponding conserved inner product on a hypersurface $\Sigma_{t}$ of constant coordinate time is given by:
+$\langle\Psi_{1}|\Psi_{2}\rangle_{FV}=\int_{\Sigma_{t}}d^{3}x\sqrt{h}\Psi_{1}^{\dagger}\eta_{FV}\Psi_{2},$ (47)
+where $h=\det(h_{ij})$ is the determinant of the induced spatial metric.
+For the FV-1/2 construction based on Eq.
+(39), this bilinear form is the natural generalization of the scalar FV charge and is conserved by the pseudo-unitary time evolution.
+When the initial data are restricted to solutions of the underlying first-order Dirac equation Eq.
+(27), the FV charge is consistent with the usual conserved Dirac current in Eq.
+(35) (up to the standard identification of the positive-frequency subspace in stationary backgrounds).
+Consequently, the Hamiltonian $\mathcal{H}$ derived in this framework is not Hermitian in the standard sense ($H^{\dagger}\neq H$) but rather pseudo-Hermitian with respect to $\eta_{FV}$.
+This condition is expressed as:
+$\mathcal{H}^{\dagger}=\eta_{FV}\mathcal{H}\eta_{FV}^{-1},$
+which ensures that the time evolution is pseudo-unitary, preserving the generalized inner product (the charge $Q$) rather than the standard probability norm Mostafazadeh (2002); Silenko (2008).
+Equivalence with the Dirac equation and elimination of spurious solutions
+Because the FV Hamiltonian is obtained from the squared Dirac operator, it is worth recalling a standard subtlety: every solution of the Dirac equation is a solution of the squared equation, but the converse is not true in general.
+The second-order equation admits additional (spurious) solutions that do not satisfy the original first-order constraint.
+In the present work, the FV formulation is derived starting from Eq.
+(27), and we therefore regard the FV Hamiltonian evolution as a Hamiltonian representation of Dirac dynamics on the physical subspace.
+Concretely, given an FV state $\Psi_{FV}=(\phi,\chi)^{T}$, the underlying Dirac spinor is reconstructed as $\Psi=\phi+\chi$, while the auxiliary combination $\phi-\chi$ is fixed by the first-order relation in Eq.
+(45).
+If the initial data obey the Dirac constraint, then the FV evolution preserves this constraint and no spurious solutions are generated.
+This “Dirac restriction” is the precise sense in which the FV-1/2 formalism is equivalent to the standard covariant Dirac theory.
+Using the FV transformation, the Klein–Gordon-type equation obtained by squaring the Dirac operator is rewritten as a first-order equation for $\psi$.
 Substituting the definitions of $\phi$ and $\chi$, we derive a coupled system of equations:
@@
 where the Hamiltonian operator $\mathcal{H}$ in terms of the components becomes:
-$\mathcal{H}=\begin{pmatrix}\mathrm{Y}+\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}&-\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)\\ -\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)&-\mathrm{Y}-\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}\end{pmatrix}.$ (22)
+$\mathcal{H}=\begin{pmatrix}\mathrm{Y}+\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}&-\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)\\ -\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)&-\mathrm{Y}-\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}\end{pmatrix}$.
+(49)
 For practical cases, simplifications often involve: (i) Assuming specific forms of the metric $g_{\mu\nu}$, (ii) Neglecting terms like $\frac{R}{4}$ or $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ in weak field approximations, and (iii) Approximating $\mathrm{Y}$ in terms of first-order derivatives of the metric and connections.
@@
 which accounts for the kinetic terms and electromagnetic interactions.
-III.1 (1+2) dimension
+IV.1 (1+2) dimension
 The term $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ represents the interaction between the electromagnetic field tensor $F_{\mu\nu}$ and the spinor field via the spin matrix $\sigma^{\mu\nu}$.
@@
 $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e(\gamma^{0}\gamma^{1}E_{x}+\gamma^{0}\gamma^{2}E_{y}).$ (60)
-III.2 in (1+3)-dimensions
+IV.2 in (1+3)-dimensions
 To evaluate $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ in $(1+3)$-dimensional spacetime, we need to compute the interaction between the electromagnetic field tensor $F_{\mu\nu}$ and the spin matrices $\sigma^{\mu\nu}.$
@@
 The magnetic field term is:
-$-e\vec{\Sigma}\cdot\vec{B}=-e\begin{pmatrix}\vec{\sigma}&0\\ 0&\vec{\sigma}\end{pmatrix}\cdot\vec{B}.$ (43)
+$e\vec{\Sigma}\cdot\vec{B}=-e\begin{pmatrix}\vec{\sigma}&0\\ 0&\vec{\sigma}\end{pmatrix}\cdot\vec{B}.$ (70)
 Combining both terms:
 $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}(\gamma^{1}E_{x}+\gamma^{2}E_{y}+\gamma^{3}E_{z})-e\begin{pmatrix}\vec{\sigma}\cdot\vec{B}&0\\ 0&\vec{\sigma}\cdot\vec{B}\end{pmatrix}$. (71)
-III.2.1 Special Case: Uniform Fields
-  • Uniform Magnetic Field $(\vec{E}=0)$
-    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=-e\vec{\Sigma}\cdot\vec{B}.$ (45)
-  • Uniform Electric Field $(\vec{B}=0)$
-    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}.$ (46)
-IV Conclusion
-The Feshbach-Villars formalism for a spin-1/2 particle in curved spacetime provides a robust method for exploring relativistic quantum mechanics in the presence of gravitational and electromagnetic fields.
-By deriving the Hamiltonian form of the Dirac equation and applying the FV transformation, we have developed a two-component formulation that clearly separates positive and negative energy states.
-This formalism presents several advantages:
-  • It allows a more intuitive interpretation of the wavefunction components with respect to particle and antiparticle contributions.
-  • The resulting Hamiltonian, expressed in terms of Pauli matrices, facilitates the analysis of spin-dependent effects in curved spacetime.
-  • The approach naturally incorporates both gravitational and electromagnetic interactions, making it suitable for exploring various physical scenarios.
-Our analysis in both (1+2) and (1+3) dimensions showcases the versatility of the formalism and highlights the specific forms of spin-field interactions across different spacetime configurations.
-The explicit treatment of the term involving the electromagnetic field tensor and spin matrices provides a clear understanding of how spin couples to external fields in curved spacetime.
-This work lays the groundwork for future studies on relativistic quantum effects in strong gravitational fields, the behavior of spin in curved spacetime, and potential applications in areas such as early universe cosmology and the physics of compact objects.
-Further research could investigate the implications of this formalism for quantum field theory in curved spacetime and explore possible experimental tests of the predicted effects.
-?refname?
-  • 1 (2025) H. Feshbach and F. M. H. Villars, Rev. Modern Phys. 30, 24 (1958).
-  • 2 (2025) B. A. Robson and D. S. Staudte, J. Phys. A : Math. Gen 29, 157 (1996).
-  • 3 (2025) D. S. Staudte, J. Phys. A 29, 169 (1996).
-  • 4 (2025) M. Merad, L. Chetouani, and A. Bounames, Phys. Lett. A 267, 225 (2000).
-  • 5 (2025) A. Bounames and L. Chetouani, Phys. Lett. A 279, 139 (2001).
-  • 6 (2025) S. Haouat and L. Chetouani, Eur. Phys. J. C 41, 297 (2005).
-  • 7 (2025) B Guettou1 and L Chetouani, Phys. Scr. 74, 12–18 (2006).
-  • 8 (2025) N. Brown, Z. Papp, and R. M. Woodhouse, Few-Body Systems 57, 103 (2015).
-  • 9 (2025) B. Motamedi, T. Shannon, and Z. Papp, Few-Body Systems 60, 65 (2019).
-  • 10 (2025) D. Wingard, A. Garcia Vallejo and Z. Papp, Few-Body Syst 65, 30 (2024).
-  • 11 (2025) A. J. Silenko, Phys. Rev. A 77, 012116 (2008).
-  • 12 (2025) A. J. Silenko, Phys. Rev. D 88, 045004 (2013).
-  • 13 (2025) A. Bouzenada, A. Boumali, E.O. Silva, Ann. Phys. 458, 169479 (2023)
-  • 14 (2025) A. Bouzenada, A. Boumali, R. Vitória, F. Ahmed, M. Al-Raeei, Nucl. Phys. B 994, 116288 (2023)
-  • 15 (2025) A. Bouzenada, A. Boumali, F. Ahmed, Nucl. Phys. B 1007, 116682 (2024)
-  • 16 (2025) A.Bouzenada and A.Boumali. Ann. Physics 27, 169302 (2023).
-  • 17 (2025) E. Schrodinger, Sitz. Preuss. Akad. Wiss. Berlin 105 (1932).
+Now, to simplify the handling of our quantum mechanics problem, let us work in a space of lower dimensions, and because there is no structure in the $z$ direction, we can remove it.
+Thus, the metric of a static cosmic string with cylindrical symmetry in two dimensions has the following form Vilenkin (1981); Vilenkin and Shellard (1994)
+$ds^{2}=dt^{2}-dr^{2}-\alpha^{2}r^{2}d\varphi^{2}$ (72)
+where the components of the metric and the inverse metric tensors are, respectively,
+$g_{\mu\nu}=\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&-(\alpha r)^{2}\end{array}\right),\quad g^{\mu\nu}=\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&\frac{-1}{(\alpha r)^{2}}\end{array}\right)$ (73)
+Note here that, in the Hamiltonian formulation, the evolution parameter is the coordinate time $t$ that labels the foliation by spacelike hypersurfaces $\Sigma_{t}$ defined by $t=\text{const}$.
+This is the same parameter that appears in the Schrödinger-type equation $i\partial_{t}\Psi=\mathcal{H}\Psi$.
+For a general \$3+1\$ decomposition of the spacetime metric (ADM formalism), the line element is written as:
+$ds^{2}=(N^{2}-h_{ij}N^{i}N^{j})dt^{2}-2h_{ij}N^{j}dx^{i}dt-h_{ij}dx^{i}dx^{j},$ (74)
+where $N(t,x)$ is the lapse function and $N^{i}(t,x)$ is the shift vector Arnowitt1962; Wald (1984); Gourgoulhon (2007).
+Physically, the lapse function $N$ controls the proper-time separation between neighboring hypersurfaces as measured by Eulerian observers normal to $\Sigma_{t}$.
+The shift vector $N^{i}$ encodes how the spatial coordinates are shifted when evolving from $\Sigma_{t}$ to $\Sigma_{t+dt}$, representing the gauge freedom in the choice of spatial coordinates.
+Consequently, the Hamiltonian flow is naturally parametrized by $t$.
+The relationship between $t$ and the proper time $\tau$ along a given worldline depends on $N$ and $N^{i}$.
+For the static diagonal cosmic-string metric considered in Section V, we have $g_{0i}=0$ and $g_{00}=1$.
+This implies a lapse of $N=\sqrt{g_{00}}=1$ and a vanishing shift vector $N^{i}=0$.
+Therefore, for static observers (fixed spatial coordinates), the proper time satisfies $d\tau=\sqrt{g_{00}}dt=dt$.
+In this specific case, the Hamiltonian-flow parameter coincides with both the metric time coordinate and the proper time of static observers.
+Now, to evaluate the operators $Y$ and $O$ in the Feshbach-Villars (FV) formalism for a spin-1/2 particle in the static cosmic string spacetime, we start with the metric:
+$ds^{2}=dt^{2}-dr^{2}-\alpha^{2}r^{2}d\varphi^{2},$ (75)
+with the metric tensor $g_{\mu\nu}=\text{diag}(1,-1,-\alpha^{2}r^{2})$. The inverse metric is $g^{\mu\nu}=\text{diag}(1,-1,-1/(\alpha^{2}r^{2}))$, and $\sqrt{-g}=\alpha r$.
+From the definition in the FV formalism:
+$Y=\frac{1}{2g_{00}\sqrt{-g}}\{D_{i},\sqrt{-g}g^{0i}\}.$ (76)
+Since the metric is diagonal, $g^{0i}=0$ for all spatial indices $i$. This implies:$Y=0.$ The operator $O$ is given by:
+$O=\frac{1}{2m}\left[\frac{1}{\sqrt{-g}}D_{i}\left(\sqrt{-g}g^{ij}D_{j}\right)+\frac{ie}{2}\frac{F_{\mu\nu}\sigma^{\mu\nu}}{g_{00}}\right].$ (77)
+Assuming no electromagnetic field ($F_{\mu\nu}=0$) and focusing on the gravitational contribution, we simplify:
+$O=\frac{1}{2m}\frac{1}{\sqrt{-g}}D_{i}\left(\sqrt{-g}g^{ij}D_{j}\right).$ (78)
+In cylindrical coordinates ($r,\varphi$):
+$O=\frac{1}{\alpha r}\left[-D_{r}(\alpha rD_{r})-D_{\varphi}\left(\frac{1}{\alpha r}D_{\varphi}\right)\right].$ (79)
+Expanding the covariant derivatives ($D_{i}=\partial_{i}+\Gamma_{i}$) and simplifying for the static cosmic string metric (where the spin connection $\Gamma_{i}$ contributes terms proportional to $\alpha$), we obtain:
+$O=-\frac{1}{2m}\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right].$ (80)
+The FV Hamiltonian for the spin-1/2 particle becomes:
+$H=\sigma_{z}\frac{m}{2}+\sigma_{x}O,$ (81)
+leading to the coupled system:
+$i\partial_{t}\begin{pmatrix}\phi\\ \chi\end{pmatrix}=\left(\frac{m}{2}\sigma_{z}-\frac{1}{2m}\sigma_{x}\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right]\right)\begin{pmatrix}\phi\\ \chi\end{pmatrix}$.
+(82)
+Fron the last equation, The system of equations for $\phi$ and $\chi$ derived from the given Hamiltonian is:
+$i\partial_{t}\phi=\dfrac{m}{2}\phi-\dfrac{1}{2m}\left(\partial_{r}^{2}+\dfrac{1}{r}\partial_{r}+\dfrac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right)\chi$ (83)
+$i\partial_{t}\chi=-\dfrac{m}{2}\chi-\dfrac{1}{2m}\left(\partial_{r}^{2}+\dfrac{1}{r}\partial_{r}+\dfrac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}\right)\phi$ (84)
+The FV formalism rewrites the wave function as:
+$\psi=\phi+\chi$ (85)
+and relates the time derivative to the difference:
+$i\partial_{t}\psi=m(\phi-\chi)$ (86)
+This allows you to express $\phi$ and $\chi$ in terms of $\psi$ and its time derivative:
+$\phi=\frac{1}{2}\left(\psi+\frac{i}{m}\partial_{t}\psi\right)$ (87)
+$\chi=\frac{1}{2}\left(\psi-\frac{i}{m}\partial_{t}\psi\right)$ (88)
+From some algebraic steps, we obtain a differential equation as:
+$\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}+E^{2}-m^{2}\right]\psi=0$ (89)
+To solve the partial differential equation, we employ separation of variables in cylindrical coordinates. Assuming
+$\psi(r,\varphi)=R(r)\Phi(\varphi)$ (90)
+, the equation splits into angular and radial components. The angular part satisfies:
+$\Phi^{\prime\prime}(\varphi)+\alpha^{2}l^{2}\Phi(\varphi)=0,$ (91)
+with periodic boundary conditions $\Phi(\varphi+2\pi)=\Phi(\varphi)$. The solution is:
+$\Phi(\varphi)=e^{il\varphi},\quad l\in\mathbb{Z}.$ (92)
+Substituting $\Phi(\varphi)$ into the original equation yields:
+$R^{\prime\prime}(r)+\frac{1}{r}R^{\prime}(r)+\left(k^{2}-\frac{l^{2}}{\alpha^{2}r^{2}}\right)R(r)=0,$ (93)
+where $k^{2}=E^{2}-m^{2}$.
+This is a Bessel equation of order $\nu=l/\alpha$. The general solution is Andrews et al. (1999); Arfken and Weber (2005):
+$R(r)=AJ_{l/\alpha}(kr)+BY_{l/\alpha}(kr),$ (94)
+where $J_{\nu}$ and $Y_{\nu}$ are Bessel functions of the first and second kind. For regularity at $r=0$, $B=0$, leaving:
+$R(r)=AJ_{l/\alpha}(kr).$ (95)
+Combining the radial and angular components, the general solution is:
+$\psi(r,\varphi)=\sum_{l=-\infty}^{\infty}A_{l}J_{l/\alpha}(kr)e^{il\varphi}.$ (96)
+The energy $E$ is related to the wave number $k$ by:
+$E=\pm\sqrt{m^{2}+k^{2}}.$ (97)
+Boundary conditions (e.g., $R(R)=0$ at radius $R$) quantize $k$ via zeros of $J_{l/\alpha}(kR)$.
+IV.2.1 FV-1/2 oscillator
+Now, to include the Dirac oscillator term $im\omega\beta r$ into eq.
+(2), we proceed with the following substitution in the radial momentum component $\partial_{r}\rightarrow\partial_{r}+m\omega\rho$.
+Hence, eq.
+(89) is transformed into
+$\left[\left(\partial_{r}-m\omega\rho\right)\left(\partial_{r}+m\omega\rho\right)+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}+E^{2}-m^{2}\right]\psi=0$ (98)
+or
+$\left[\partial_{r}^{2}-m^{2}\omega^{2}\rho^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}+E^{2}-m^{2}+m\omega\right]\psi=0$ (99)
+To solve the given partial differential equation (PDE) we employ the separation of variables method, assuming a solution of the form
+$\psi(r,\varphi)=R(r)\Phi(\varphi)$ (100)
+Separate the angular component by substituting $\psi=R(r)\Phi(\varphi)$ into the PDE and dividing by $R\Phi$:
+$\frac{1}{\alpha^{2}r^{2}}\frac{\partial_{\varphi}^{2}\Phi}{\Phi}=-\nu^{2}\quad\Rightarrow\quad\Phi^{\prime\prime}+\alpha^{2}\nu^{2}\Phi=0.$ (101)
+The periodic boundary condition $\Phi(\varphi+2\pi)=\Phi(\varphi)$ quantizes $\nu$ as an integer ($\nu\in\mathbb{Z}$). The solution is:
+$\Phi(\varphi)=e^{i\nu\alpha\varphi}.$ (102)
+Substitute $\Phi(\varphi)$ back into the PDE to obtain the radial equation:
+$R^{\prime\prime}+\frac{1}{r}R^{\prime}+\left[E^{2}-m^{2}+m\omega-m^{2}\omega^{2}r^{2}-\frac{\nu^{2}}{\alpha^{2}r^{2}}\right]R=0.$ (103)
+This resembles a quantum harmonic oscillator equation with an additional $1/r^{2}$ term Greiner (2001); Flugge (1974) . The Analyze Asymptotic Behavior, we have
+  • At $r\to 0$: The dominant term is $\frac{\nu^{2}}{\alpha^{2}r^{2}}$.
+  • At $r\to\infty$: The dominant term is $m^{2}\omega^{2}r^{2}$.
+This suggests an ansatz:
+$R(r)=r^{|\nu|/\alpha}e^{-\frac{1}{2}m\omega r^{2}}F(r)$ (104)
+The power of $r$ handles the singularity at the origin, and the exponential ensures normalizability at infinity.
+Plug $R(r)$ into the radial equation and simplify.
+You will obtain a differential equation for $F(r)$ of the form:
+$F^{\prime\prime}(r)+\left(\frac{2|\nu|/\alpha+1}{r}-2m\omega r\right)F^{\prime}(r)+\left[\kappa-2m\omega\frac{|\nu|}{\alpha}-m\omega\right]F(r)=0,$ (105)
+where
+$\kappa=E^{2}-m^{2}+m\omega$ (106)
+we recognize it as a form of the associated Laguerre equation Greiner (2001); Flugge (1974) .
+Let $x=m\omega r^{2}$. Changing variables and simplifying reduces the equation to:
+$x\frac{d^{2}F}{dx^{2}}+\left(\frac{|\nu|}{\alpha}+1-x\right)\frac{dF}{dx}+\left[\frac{\kappa-2m\omega|\nu|/\alpha-m\omega}{4m\omega}\right]F=0.$ (107)
+This matches the standard associated Laguerre equation:
+$xy^{\prime\prime}+(k+1-x)y^{\prime}+ny=0,$ (108)
+where: $k=|\nu|/\alpha$, and
+$n=\frac{\kappa-2m\omega|\nu|/\alpha-m\omega}{4m\omega}$ (109)
+For polynomial solutions (associated Laguerre polynomials $L_{n}^{(k)}(x)$), $n$ must be a non-negative integer.
+This imposes:
+$\kappa=4m\omega n+2m\omega\frac{|\nu|}{\alpha}+m\omega.$ (110)
+Recalling $\kappa=E^{2}-m^{2}+m\omega$, we solve for $E$:
+$E^{2}=m^{2}+4m\omega n+2m\omega\frac{|\nu|}{\alpha}.$ (111)
+The discrete energy levels are:
+$E=\pm\sqrt{m^{2}+2m\omega\left(2n+\frac{|\nu|}{\alpha}\right)},$ (112)
+where $n=0,1,2,\ldots$ is the radial quantum number, and $\nu\in\mathbb{Z}$ is the angular quantum number.
+The full solution combines radial and angular components:
+$\psi(r,\varphi)=\sum_{n,\nu}C_{n,\nu}\,r^{|\nu|/\alpha}e^{-m\omega r^{2}/2}L_{n}^{|\nu|/\alpha}(m\omega r^{2})\,e^{i\nu\alpha\varphi},$ (113)
+where $C_{n,\nu}$ are normalization constants. The energy spectrum is discrete and depends on $n$, $\nu$, and $\alpha$.
+V FV oscillator in a spinning cosmic string with torsion and dislocation
+In this third example we consider a spinning cosmic string carrying both a time–like torsion (frame–dragging) and a space–like torsion (screw dislocation).
+A convenient form of the corresponding line element (in cylindrical coordinates) is Puntigam and Soleng (1997); Ozdemir (2005); Jusufi (2016)
+$ds^{2}=\left(dt+a\,d\varphi\right)^{2}-dr^{2}-\alpha^{2}r^{2}d\varphi^{2}-\left(dz+\beta\,d\varphi\right)^{2},$ (114)
+where $\alpha\in(0,1]$ is the angular–deficit parameter, $a$ is proportional to the linear density of angular momentum (time–like torsion) and $\beta$ encodes the screw–dislocation (space–like torsion).
+Note that $g_{\varphi\varphi}=a^{2}-\alpha^{2}r^{2}-\beta^{2}$; in the regime $a^{2}>\beta^{2}$ one must restrict the radial coordinate to $r>r_{0}=\sqrt{a^{2}-\beta^{2}}/\alpha$ in order to avoid unphysical closed time–like curves, so that $r=r_{0}$ plays the role of a geometric hard wall Garah and Boumali (2025b).
+A convenient orthonormal coframe $\{\hat{\theta}^{a}\}$ adapted to the line element (114) is
+$\hat{\theta}^{0}=dt+a\,d\varphi,\qquad\hat{\theta}^{1}=dr,\qquad\hat{\theta}^{2}=\alpha r\,d\varphi,\qquad\hat{\theta}^{3}=dz+\beta\,d\varphi,$ (115)
+so that $ds^{2}=\eta_{ab}\,\hat{\theta}^{a}\hat{\theta}^{b}$ with $\eta_{ab}=\mathrm{diag}(1,-1,-1,-1)$.
+The corresponding tetrad $e^{a}{}_{\mu}$ and its inverse $e_{a}{}^{\mu}$ (defined by $\hat{\theta}^{a}=e^{a}{}_{\mu}\,dx^{\mu}$ and $dx^{\mu}=e_{a}{}^{\mu}\,\hat{\theta}^{a}$) read
+$\left(e^{a}{}_{\mu}\right)=\begin{pmatrix}1&0&a&0\\ 0&1&0&0\\ 0&0&\alpha r&0\\ 0&0&\beta&1\end{pmatrix},\qquad\left(e_{a}{}^{\mu}\right)=\begin{pmatrix}1&0&-\dfrac{a}{\alpha r}&0\\ 0&1&0&0\\ 0&0&\dfrac{1}{\alpha r}&0\\ 0&0&-\dfrac{\beta}{\alpha r}&1\end{pmatrix}$, (116)
+where the coordinate ordering is $(t,r,\varphi,z)$ and the local Lorentz index ordering is $(0,1,2,3)$.
+In the notation of Sec.
+II, the curved-space Dirac matrices are $\gamma^{\mu}=e_{a}{}^{\mu}\gamma^{a}$, where $\gamma^{a}$ satisfy $\{\gamma^{a},\gamma^{b}\}=2\eta^{ab}$.
+In the standard Dirac representation,
+$\gamma^{0}=\begin{pmatrix}\mathbb{I}_{2}&0\\ 0&-\mathbb{I}_{2}\end{pmatrix},\qquad\gamma^{i}=\begin{pmatrix}0&\sigma^{i}\\ -\sigma^{i}&0\end{pmatrix}\ (i=1,2,3),\qquad\alpha^{i}=\gamma^{0}\gamma^{i},\ \beta=\gamma^{0},$ (117)
+with $\sigma^{i}$ the Pauli matrices and $\mathbb{I}_{2}$ the $2\times 2$ identity. Using (116), the explicit curved matrices needed in the Hamiltonian/FV form are
+$\gamma^{t}=\gamma^{0}-\frac{a}{\alpha r}\gamma^{2},\qquad\gamma^{r}=\gamma^{1},\qquad\gamma^{\varphi}=\frac{1}{\alpha r}\gamma^{2},\qquad\gamma^{z}=\gamma^{3}-\frac{\beta}{\alpha r}\gamma^{2}.$ (118)
+For $r\neq 0$, the only $r$-dependent coframe component is $\hat{\theta}^{2}=\alpha r\,d\varphi$, and the torsion is distributional at the string core (screw dislocation) while vanishing in the exterior region.
+In the exterior region one finds the single nonvanishing Lorentz connection component
+$\omega_{\varphi}{}^{12}=1-\alpha,\qquad\Rightarrow\qquad\Gamma_{\varphi}=\frac{1}{4}\,\omega_{\varphi}{}^{12}\,\gamma_{12}=\frac{1-\alpha}{4}\,\gamma^{1}\gamma^{2},$ (119)
+where $\Gamma_{\mu}=\frac{1}{4}\omega_{\mu}{}^{ab}\gamma_{ab}$ is the spinor connection used in Sec.
+II and $\gamma_{ab}=\frac{1}{2}[\gamma_{a},\gamma_{b}]$.
+The parameter $\alpha\in(0,1]$ encodes the angular deficit (for a straight string $\alpha=1-4\mu$), $a$ is proportional to the angular momentum per unit length (“spinning” string), and $\beta$ is the screw-dislocation/torsion parameter (Burgers vector per $2\pi$) that shifts the identification along $z$ under $\varphi\to\varphi+2\pi$.
+In the FV–oscillator problem, $m$ and $\omega$ denote the particle mass and the oscillator frequency, while $l\in\mathbb{Z}$ and $k\in\mathbb{R}$ are the azimuthal and longitudinal quantum numbers, respectively.
+Because the background is stationary and invariant under translations along $z$, we separate variables as
+$\psi\left(t,r,\varphi,z\right)=e^{-iEt}\,e^{il\varphi}\,e^{ikz}\,R(r),$ (120)
+with $l\in\mathbb{Z}$ and $k\in\mathbb{R}$.
+The mixed components $g_{t\varphi}$ and $g_{z\varphi}$ imply that the azimuthal dependence enters through the combination $(\partial_{\varphi}+a\,\partial_{t}-\beta\,\partial_{z})$ .
+Consequently, the effective angular index is shifted to
+$\nu_{\mathrm{eff}}=l+aE-\beta k.$ (121)
+Following the same FV–oscillator steps as in the previous subsection, we obtain the radial equation
+$\frac{d^{2}R}{dr^{2}}+\frac{1}{r}\frac{dR}{dr}+\left[\left(E^{2}-m^{2}-k^{2}\right)-m^{2}\omega^{2}r^{2}-\frac{\nu_{\mathrm{eff}}^{2}}{\alpha^{2}r^{2}}\right]R=0.$ (122)
+Upon introducing the variable $x=m\omega r^{2}$, the regularity at the origin and normalizability yield the quantization condition
+$E^{2}=m^{2}+k^{2}+4m\omega n+\frac{2m\omega}{\alpha}\left|\nu_{\mathrm{eff}}\right|,\qquad n=0,1,2,\ldots$ (123)
+which is implicit in the energy through $\nu_{\mathrm{eff}}$ (a characteristic feature of spinning backgrounds).
+For the branch $\nu_{\mathrm{eff}}>0$, Eq.
+(123) becomes a quadratic equation for $E$ and can be written explicitly as
+$E=\frac{m\omega a}{\alpha}\pm\sqrt{\left(\frac{m\omega a}{\alpha}\right)^{2}+m^{2}+k^{2}+4m\omega n+\frac{2m\omega}{\alpha}\left(l-\beta k\right)}.$ (124)
+For $\nu_{\mathrm{eff}}<0$ we obtain instead
+$E=-\frac{m\omega a}{\alpha}\pm\sqrt{\left(\frac{m\omega a}{\alpha}\right)^{2}+m^{2}+k^{2}+4m\omega n-\frac{2m\omega}{\alpha}\left(l-\beta k\right)}.$ (125)
+In the limits $a\rightarrow 0$ and $\beta\rightarrow 0$ we recover the static cosmic string result of Eq.
+(111); the flat–space Moshinsky spectrum follows further for $\alpha\rightarrow 1$.
+This provides a nontrivial consistency check of the FV construction: in all limits where the spectrum is known from the standard Dirac analysis, the observable eigenvalues coincide, as expected for a Hamiltonian reformulation.
+VI Conclusion
+This work extends the Feshbach–Villars (FV) formalism to spin–$1/2$ particles in curved spacetime by deriving a Hamiltonian representation of the covariant Dirac dynamics and implementing a diagonal FV transformation.
+The resulting two-component framework makes the separation between positive- and negative-frequency sectors explicit and provides a compact way to track gravitational and electromagnetic couplings, including spin–field interaction terms, in both $(1+2)$ and $(1+3)$ dimensions.
+Discussion.
+The applications to static cosmic-string geometries (with and without oscillator couplings) illustrate how geometric parameters (such as the deficit angle and torsion/dislocation data) control the quantization conditions, spectra, and wave-function profiles.
+In this sense, the FV representation serves as a practical bridge between standard Dirac dynamics and second-order formulations in which particle–antiparticle structure is less transparent.
+In addition, we have made explicit two points that are sometimes only implicit in second-order approaches: (i) the FV Hamiltonian is pseudo-Hermitian with respect to a simple metric operator $\eta_{FV}=\sigma_{3}\otimes\mathbb{I}_{4}$, so that the conserved quantity is a charge-type norm, and (ii) the equivalence with the standard Dirac equation is understood as a restriction to the physical (Dirac) subspace that removes the spurious solutions of the squared equation.
+Finally, the role of stationarity is particularly transparent: when $g_{0i}\neq 0$ (nonvanishing ADM shift), the FV operator $\mathrm{Y}$ does not vanish and this directly feeds into the energy quantization.
+The spinning string example exhibits this through the effective angular index $\nu_{\mathrm{eff}}=l+aE-\beta k$ and the analytic spectra in Eqs.
+(124)–(125).
+Limitations.
+The present approach is a single-particle relativistic quantum-mechanical description on a prescribed background.
+It does not include intrinsically field-theoretic effects in curved spacetime, such as particle creation, vacuum polarization, or gravitational backreaction, and it does not resolve the well-known ambiguity of particle concepts in generic nonstationary geometries.
+Moreover, global issues (e.g., the existence of a suitable foliation and boundary conditions ensuring self-adjointness in the FV inner product) can restrict the applicability of a Hamiltonian treatment.
+These limitations help explain why semiclassical and quantum-field-theoretic methods have long been central to the study of curved-spacetime phenomena.
+Within its domain of validity, the FV formulation developed here provides a systematic platform for further studies, including extensions to broader classes of stationary metrics, controlled comparisons with semiclassical/QFT treatments, and investigations of spinor dynamics in other topologically nontrivial backgrounds.
+VII Data Availability Statement
+No Data associated in the manuscript
 Appendix A Proof of Eq. (39)
@@
 The gamma matrices satisfy the Clifford algebra:
-$\gamma^{\mu}\gamma^{\nu}=g^{\mu\nu}+\sigma^{\mu\nu},\quad\text{where }\sigma^{\mu\nu}=\frac{1}{2}\left[\gamma^{\mu},\gamma^{\nu}\right].$ (51)
+[…]
 Substitute:
-$-\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=-g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}-\sigma^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (52)
+$\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=-g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}-\sigma^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (131)
 The covariant derivative $\mathcal{D}_{\mu}=\partial_{\mu}+\Gamma_{\mu}+\text{spin connection}$ satisfies:
-$[\mathcal{D}_{\mu},\mathcal{D}_{\nu}]=\frac{1}{4}R_{\mu\nu\lambda\sigma}\gamma^{\lambda\sigma}-iqF_{\mu\nu}.$ (53)
+$\mathcal{D}_{\mu},\mathcal{D}_{\nu}]=\frac{1}{4}R_{\mu\nu\lambda\sigma}\gamma^{\lambda\sigma}-iqF_{\mu\nu}.$ (132)
 From this:
@@
 $\left(i\gamma^{\mu}\mathcal{D}_{\mu}\right)^{2}=-\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}\right).$ (135)
-Adding the mass term $-m^{2}$:
+Adding the mass term $m^{2}$:
 $\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\left[i\gamma^{\mu}\mathcal{D}_{\mu}+m\right]\Psi=\left(-\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)+\frac{1}{4}R-\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0.$ (136)
@@
 Here, the derivation is exact under the assumptions of: Minimal coupling to gravity and electromagnetism, Metric compatibility $(\nabla_{\mu}g_{\alpha\beta}=0)$ and Torsion-free spacetime.
-It is important to mention that (58) was originally derived by Schrodinger (17, 2025).
+It is important to mention that (137) was originally derived by Schrodinger Schrodinger1932.
+Appendix B Proof of Eq. (89)
+The FV formalism rewrites the wave function as:
+$\psi=\phi+\chi$ (138)
+and relates the time derivative to the difference:
+$i\partial_{t}\psi=m(\phi-\chi)$ (139)
+This allows you to express $\phi$ and $\chi$ in terms of $\psi$ and its time derivative:
+$\phi=\frac{1}{2}\left(\psi+\frac{i}{m}\partial_{t}\psi\right)$ (140)
+$\chi=\frac{1}{2}\left(\psi-\frac{i}{m}\partial_{t}\psi\right)$ (141)
+Given:
+$i\partial_{t}\phi=\frac{m}{2}\phi-\frac{1}{2m}\mathcal{D}\chi$ (142)
+$i\partial_{t}\chi=-\frac{m}{2}\chi-\frac{1}{2m}\mathcal{D}\phi$ (143)
+where
+$\mathcal{D}=\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}$ (144)
+Differentiate the first equation with respect to time and use the second to eliminate $\chi$:
+  • Start with the first equation:
+    $i\partial_{t}\phi=\frac{m}{2}\phi-\frac{1}{2m}\mathcal{D}\chi$ (145)
+  • Take the time derivative:
+    $i\partial_{t}(i\partial_{t}\phi)=i\partial_{t}\left(\frac{m}{2}\phi-\frac{1}{2m}\mathcal{D}\chi\right)$ (146)
+    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi-\frac{1}{2m}\mathcal{D}i\partial_{t}\chi$ (147)
+  • Substitute $i\partial_{t}\chi$ from the second equation:
+    $i\partial_{t}\chi=-\frac{m}{2}\chi-\frac{1}{2m}\mathcal{D}\phi$ (148)
+  • Plug into the previous expression:
+    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi-\frac{1}{2m}\mathcal{D}\left(-\frac{m}{2}\chi-\frac{1}{2m}\mathcal{D}\phi\right)$ (149)
+    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi+\frac{m}{2}\frac{1}{2m}\mathcal{D}\chi+\frac{1}{4m^{2}}\mathcal{D}^{2}\phi$ (150)
+    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi+\frac{1}{2}\mathcal{D}\chi+\frac{1}{4m^{2}}\mathcal{D}^{2}\phi$ (151)
+Now, recall $\psi=\phi+\chi$. Using the above and similar steps for $\chi$, you can combine to get a second-order equation for $\psi$.
+Following the FV formalism for spin-1/2 in curved spacetime, the second-order equation for $\psi$ in your geometry (static cosmic string, 2D) is:
+$\left[\partial_{t}^{2}-\mathcal{D}+m^{2}\right]\psi=0$ (152)
+where
+$\mathcal{D}=\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}$ (153)
+This matches the generalized Klein-Gordon equation obtained by squaring the Dirac operator and applying the FV transformation.